OpenStudy (anonymous):

I need to see how to convert y=(x-1)/(2x+3) into x=(3y+1)/(1-2y). I know I need to cross multiply, but I get stuck with 2xy+3y=x-1 ... what is the next step?

OpenStudy (anonymous):

(2x+3)y=x-1 2xy+3y=x-1 2xy-x=-3y-1 x(2y-1)=-3y-1 x= -3y-1/(2y-1)

OpenStudy (anonymous):

(3y+1)/(-2y+1) negative cancel